A firm purchases very large quantity of metal offcuts and wishes to know the average weight of an offcut. A random sample of 625 offcuts is weighted and it is found that the mean sample weight 150 grams with a sample standard deviation of 30 grams. What is the standard error of the mean? What would be the standard error if sample size was 1,225?  

 



1
Expert's answer
2021-08-05T08:04:08-0400

Standard error "\sigma_x" for n samples with standard deviation "\sigma" is given by a formula

"\sigma_x = \frac{\sigma}{\sqrt{n}}"

So for 625 samples "\sigma_x = \frac{30}{\sqrt{625}} = \frac{30}{25} = 1.2"

and for n = 1225 "\sigma_x = \frac{30}{\sqrt{1225}} = \frac{30}{35} \approx 0.857"